# Qubit Mapping and Routing tailored to Advanced Quantum ISAs: Not as Costly as You Think

### v1

@@ 4.3. Routing in canonical form
where $w_{g}$ and $w_{d}$ are the weight factors for the gate count and depth cost components, respectively. The last term reflects the original heuristic cost $H_{\textsc{Sabre}}=\mathrm{Avg}\{\mathrm{dist}[i,j]\}_{F}+k_{E}\,\mathrm{Avg}\{\mathrm{dist}[i,j]\}_{E}$ in Sabre (li2019tackling), while used in a unified approach accompanied with the other cost components in Canopus. Specifically, $H_{\textsc{Sabre}}$ refers to the average distance between physical qubits mapped to demanded logical interactions in the front layer $F$ and the extended set $E$ (lookahead mechanism). Here we use the “differential” average distance after inserting a $\mathrm{SWAP}$ candidate, then multiply it by the $c_{\mathrm{swap}}$ to reflect on average how an inserted $\mathrm{SWAP}$ can reduce the distance between each physical qubit pair, with the negative distance-cost increment quantified in the basis. Furthermore, the decay factor in Sabre is not needed anymore, as the $c_{g}$ and $\Delta_{\mathrm{depth}}$ guide more accurate co-optimization. Therefore, each cost component in Equation 2 refers to “cost increment”, composing a unified heuristic cost function to reconcile the multifaced routing costs.
@@ 4.5. Scalability and implementation
The overall algorithm framework to implement Canopus resembles Canopus. To efficiently implement the sophisticated $\mathrm{SWAP}$ insertion mechanism in Canopus, we develop corresponding core algorithms. Algorithm 1 specifies how the demanded data structures—the last mapped layer $L$, commutative canonical gate pairs $C$ within $L$, wire duration record DD—will be updated when adding an executable 2Q gate to the routed circuit DAG. Algorithm 2 shows how the wire durations $D$ should be correctly updated when encountering a $\mathrm{SWAP}$ insertion that satisfies the canonical gate commutativity optimization opportunity. It is also essential to evaluate the total circuit cost after mapping. Notably, all the computation processes within these algorithms are based on conditional control and operations on hashed data structures, thus with $\mathcal{O}(1)$ time complexity introduced. To quantify the synthesis cost computation of a given canonical gate, the computation is to determine which convex polytope its canonical coordinate is located in, which is a highly efficient computation process with linear time complexity. Canopus will also cache canonical gate costs it has acquired to avoid repetitive computation. Consequently, the overall scalability of Canopus is on par with that of Sabre, ensuring its practical applicability to large-scale circuits.
@@ 5. Case Studies
We validate the practical advantages of Canopus through two realistic case studies: the real-machine compilation and execution of Quantum Fourier Transform (QFT) circuits on IBM’s QPU ibm_torino, and the end-to-end simulation of compiled Quantum Low-Density Parity-Check (QLDPC) stabilizer circuits to assess its impact on the logical error rate.
@@ 5.1. QFT kernel
QFT is a fundamental subroutine in many promising quantum algorithms, such as Shor’s algorithm (shor1994algorithms), quantum phase estimation (kitaev1995quantum), etc. Given the extensive research on dedicated QFT compilers (zhang2021time; jin2024optimizing; maslov2007linear), we select the state-of-the-art compiler TOQM, which specializes in QFT optimization, as our primary baseline.
@@
A key finding is that Canopus always achieves the optimal QFT routing scheme on the 1D chain topology, while TOQM does not. Provably, the minimal number of $\mathrm{SWAP}$ insertions to route an $n$-qubit QFT is $\frac{n(n-1)}{2}-2$, that is, 2 fewer than the original $\mathrm{CPhase}$ count. This results in a perfect, symmetric butterfly circuit structure, as exemplified in Figure 8(b), with minimal #Can and 2Q circuit depth. Notably, this result is indeed optimal, surpassing the manually designed scheme previously reported as optimal by maslov2007linear where 2 more $\mathrm{SWAP}$ gates are required. This optimal scheme is irrespective of the target ISA. In contrast, our experiments show that TOQM  despite claiming to realize the scheme from (maslov2007linear), fails to reproduce it and consistently yields inferior results to Canopus, as illustrated in Figure 8.
@@
We compare compilation performance for both $6$- and $12$-qubit QFT kernels on both 1D chain and 2D square topologies, with results summarized in Table 1. On the 1D chain, Canopus always produces the theoretically optimal routing result, while TOQM does not. For the small-scale qft_6 kernel on the 2D square, Canopus also achieves the optimal routing, superior to TOQM in both #Can and 2Q depth. For the large-scale qft_12 kernel, Canopus consistently outperforms TOQM in both metrics.
@@
Figure 9. QFT kernel fidelity comparison benchmarked on IBM® Quantum Platform (ibm_torino). ibm_torino is the Heron-series QPU with native gate set $\bigl\{\mathrm{CZ},\,\sqrt{\mathrm{X}},\,\mathrm{Z}(\theta)\bigr\}$.
@@
To further validate these results, we performed real-machine experiments on IBM’s ibm_torino QPU. We compiled QFT circuits of sizes $n\in\{6,8,10,12\}$ for a 1D chain topology using both Canopus and the default Qiskit compiler. Although ibm_torino has a heavy-hex topology, it contains linear chains of sufficient size for these benchmarks. Fidelity was measured using the Hellinger fidelity between the experimental and ideal output distributions, with the number of shots set to $\textsc{max}\{4096,2^{n}\times 10\}$. A layer of Hadamard gates is appended to each circuit execution so that the ideal final state will be $\lvert 0\rangle^{\otimes n}$. In Figure 9, circuits compiled with Canopus achieve, on average, a 52.9% reduction in $\mathrm{CZ}$ gate count, a 66.4% reduction in 2Q-gate depth, and a 2.1x improvement in fidelity. These results unequivocally demonstrate the practical advantages of Canopus for QFT kernel compilation.
5.2. QLDPC stabilizer circuit
Figure 10. Logical error rate of QLDPC stabilizer circuits compiled for square (top) and heavy-hex (bottom) topologies. The y-axis shows the relative logical error rate normalized by the error rate of an ideal baseline that assumes all-to-all connectivity.
@@ 5.2. QLDPC stabilizer circuit
QLDPC codes are rapidly moving from a topic of theoretical interest to a cornerstone of experimental fault-tolerant quantum computing (FTQC) research, mainly because of their superior encoding efficiency (bravyi2024high; breuckmann2021quantum; Dennis_2002). However, due to their frequent long-range interactions (breuckmann2021quantum; panteleev2021degenerate), realizing QLDPC codes on superconducting processors with fixed, local connectivity is still hampered by significant routing overheads (wang2025demonstration).
Therefore, we attempt to compile the circuits automatically under two ISAs: (1) CX ISA with $\mathrm{CX}$ as the 2Q basis gate; (2) Stab ISA with both $\mathrm{CX}$ and $\mathrm{iSWAP}$ as basis gates, assumed to have an identical cost. Particularly, the Stab ISA aligns with practical hardware realities, e.g., both $\mathrm{CZ}$ and $\mathrm{iSWAP}$ can be natively supported by mainstream superconducting platforms (krantz2019quantum; wei2024native; arute2019quantum; nguyen2024programmable). We further build an end-to-end evaluation pipeline with QLDPC code circuit examples from (wang2025demonstration; panteleev2021degenerate). We simulate the routed circuits using stim (Gidney2021StimAF) under the circuit-level noise model (Dennis_2002; Acharya2022SuppressingQE). Finally, all syndromes are decoded using the BPOSD decoder (panteleev2021degenerate; roffe2020decoding; hillmann2024localized) to determine the logical error rate.
@@
As shown in Figure 10, Canopus consistently achieves lower logical error rates than Sabre, as the ISA-aware approach of Canopus results in compiled circuits with less $\mathrm{CX}/\mathrm{iSWAP}$ gate count and circuit depth. Under the CX, Canopus yields an average logical error suppression of 49.4% on the square topology and 11.4% on the heavy-hex topology compared to Sabre. The advantage becomes even more pronounced with the Stab, where Canopus achieves a 52.6% (square) and 29.3% (heavy-hex) error suppression. These results highlight two key findings: first, the ISA-aware mechanism in Canopus is highly effective for compiling QEC circuits, and second, the dedicated use of a hybrid $\mathrm{CX}$-$\mathrm{iSWAP}$ gate set offers a significant practical advantage for QLDPC code demonstrations on superconducting hardware.
@@ 6.1.1. ISAs and basis gate costs
We consider six different ISAs (including the conventional CX ISA) listed in Table 2. These cover a wide range of powerful basis gates from $\mathrm{CX}$-family and $\mathrm{iSWAP}$-family gates. Particularly, SQiSW (huang2023quantum) proves to a powerful ISA option and has been adopted by recent software projects (mckinney2024mirage; cirqSQiSWDecomposer). ZZPhase ISA containing three fractional $\mathrm{ZZ}(\theta)$ rotation gates (equivalently, $\left\{\sqrt[3]{\mathrm{CX}},\sqrt{\mathrm{CX}},\mathrm{CX}\right\}$) is adopted by Qiskit’s latest synthesis functionalities (peterson2022optimal; qiskitXXDecomposer). For ZZPhase and SQiSW, we also consider the mirror-enhanced version by incorporating the mirrored basis gates (cross2019validating; mckinney2024mirage) into the ISAs. We also include the Het ISA that is the composition of ZZPhase and SQiSW.
@@
(6) $\left\{\begin{array}[]{c}\mathrm{CX}:1,\,\mathrm{ZZ}(\pi/t):2/t,\,\mathrm{\sqrt{iSWAP}}:0.75,\\ \mathrm{iSWAP}:1.5,\,\mathrm{ECP}:1.25,\,\mathrm{pSWAP}(\pi/t):2-1/t\end{array}\right\},$
where $\mathrm{CX}$ gate is the unit cost. Such a setting ensures the continuity of gate costs along the critical edges in the Weyl chamber. For example, $\mathrm{pSWAP}(\pi/2)$ is equivalent to $\mathrm{iSWAP}$ and they have the same cost of $1.5$. With a specific gate family, basis gates with larger canonical coefficients usually requires proportionally longer interaction time on physical qubits, which was reflected in the cost setting. Note that this setting is a comprehensive consideration for current gate schemes and hardware-implemented gate fidelities in superconducting (chen2025efficient; arute2019quantum; wei2024native; nguyen2024programmable; acharya2024quantum). It is neither limited to a specific gate scheme nor a specific hardware platform.
@@ 6.1.4. Baselines
Hyperparameters for Sabre and Canopus are of the same settings. Each performs 10 times layout procedure, within which 5-round bidirectional passes are proceeded and each pass performs 10 trials. The best result across all attempts is selected. TOQM can obtain the deterministic routing result in one go. Compiled circuits by BQSKit, although in terms of only the 2Q gates, is also random. Thus we perform 3 trials for each input case and report the best result.
@@ 6.2. Suppression of routing overhead
Table 4. Average (geometric-mean) routing overhead across different topologies and backend ISAs.
@@
The comprehensive evaluation (on geometric-mean average) results are summarized in Table 4, across the six selected ISAs and three hardware topologies (1D chain, 2D heavy-hex, 2D square). While detailed results across the $3\times 6\times 12$ cases are listed in Table 5 and Table 6 in Appendix C. Canopus results in the lowest average (geometric-mean) routing overhead for each ISA-topology combination benchmarking, significantly surpassing other methods. Specifically, the routing overhead reduction achieved by Canopus is on average 16.06% (26.44%), 34.70% (21.25%), and 19.89% (20.72%) in terms of $C_{\mathrm{count}}$ ($C_{\mathrm{depth}}$), compared to Sabre, TOQM, and BQSKit, respectively. With Canopus, those more powerful ISAs, such as ZZPhase and SQiSW ISAs, actually exhibit significantly lower routing overhead than CX. While the baseline methods do not apparently indicate that, for which the circuit cost reduction stems from ISA rebase itself.
Notably, Sabre and BQSKit focus on gate count driven optimization, TOQM specializes in lowering circuit depth, while our Canopus involves both count and depth related optimization. Among them, TOQM leads to the worst count-related routing overhead. Even for the routing overhead in terms of depth, TOQM consistently underperforms Canopus, and it only outperforms BQSKit for 2D square topology. Besides, Canopus could manage the routing overhead in a ubiquitously low level, in terms of both count and depth costs; while every baseline fails to optimize some specific circuits. For example, TOQM/BQSKit cannot manage the routing overhead for qec9 circuit; BQSKit cannot for bv circuit, even though with more expressive ISAs.

### v3

@@ 4.3. Routing in canonical form
where $w_{g}$ and $w_{d}$ are the weight factors for the gate count and depth cost components, respectively. The last term reflects the original heuristic cost
(2) $H_{\textsc{Sabre}}=\mathrm{Avg}\{\mathrm{dist}[i,j]\}_{F}+k_{E}\,\mathrm{Avg}\{\mathrm{dist}[i,j]\}_{E}$
in Sabre, while used in a unified approach accompanied by other cost components in Canopus. Specifically, $H_{\textsc{Sabre}}$ refers to the average distance between physical qubits mapped to demanded logical interactions in the front layer $F$ and the extended set $E$ (lookahead mechanism). Here we use the “differential” average distance after inserting a $\mathrm{SWAP}$ candidate, then multiply it by the $c_{\mathrm{swap}}$ to reflect on average how an inserted $\mathrm{SWAP}$ can reduce the distance between each physical qubit pair, with the negative distance-cost increment quantified in the basis. Furthermore, the decay factor in Sabre is no longer needed, as the $c_{g}$ and $\Delta_{\mathrm{depth}}$ guide more accurate count-depth co-optimization. Therefore, each cost component in Equation 1 refers to “cost increment”, composing a unified heuristic cost function to reconcile multifaced routing costs.
@@ 4.5. Scalability and implementation
The overall algorithm framework to implement Canopus resembles Canopus. To efficiently implement the sophisticated $\mathrm{SWAP}$ insertion mechanism in Canopus, we develop corresponding core algorithms. Algorithm 1 specifies how the demanded data structures—the last mapped layer $L$, commutative canonical gate pairs $C$ within $L$, wire duration record DD—will be updated when adding an executable 2Q gate to the routed circuit DAG. Algorithm 2 shows how the wire durations $D$ should be correctly updated when encountering a $\mathrm{SWAP}$ insertion that satisfies the canonical gate commutativity optimization opportunity. It is also essential to evaluate the total circuit cost after mapping. Notably, all the computation processes within these algorithms are based on conditional control and operations on hashed data structures, thus with $\mathcal{O}(1)$ time complexity introduced. The synthesis cost of a target 2Q gate is quantified by identifying the convex polytope containing its canonical coordinate, for which the computation process is highly efficient with linear time complexity. Canopus also caches canonical gate costs it has acquired to avoid repetitive computation. Consequently, the overall scalability of Canopus is on par with that of Sabre, ensuring its practical applicability to large-scale circuits.
@@ 5. Case Studies
We validate the practical advantages of Canopus through two realistic case studies: the real-machine execution of Quantum Fourier Transform (QFT) circuits on IBM’s QPU ibm_marrakesh, and the end-to-end simulation of quantum low-Density parity-Check (qLDPC) stabilizer measurement circuits to assess its impact on the logical error rate.
@@ 5.1. QFT kernel
QFT is a fundamental subroutine in many promising quantum algorithms, such as Shor’s algorithm (shor1994algorithms), quantum phase estimation (kitaev1995quantum), etc. Given the extensive research on dedicated QFT compilers (zhang2021time; jin2024optimizing; maslov2007linear), we select the state-of-the-art compiler TOQM (zhang2021time), which specializes in QFT optimization, as our primary baseline.
@@
A key finding is that Canopus always achieves the optimal QFT routing scheme on the 1D chain topology, while TOQM does not. Provably, the minimal number of $\mathrm{SWAP}$ insertions to route an $n$-qubit QFT is $\frac{n(n-1)}{2}-2$, that is, 2 fewer than the original $\mathrm{CPhase}$ count. This results in a perfect, symmetric butterfly circuit structure, as exemplified in Figure 8(b), with minimal $\mathrm{Can}$ and 2Q circuit depth. Notably, this result is indeed optimal, surpassing the manually designed scheme previously reported as optimal by Maslov (maslov2007linear) where 2 more $\mathrm{SWAP}$ gates are required. This optimal scheme is irrespective of the target ISA. In contrast, our experiments show that TOQM  despite claiming to realize the scheme from (maslov2007linear), fails to reproduce it and consistently yields inferior results to Canopus, as illustrated in Figure 8.
@@
We compare compilation performance for both $6$- and $12$-qubit QFT kernels on both 1D chain and 2D square topologies, with results summarized in Table 1. On the 1D chain, Canopus always produces the theoretically optimal routing result, while TOQM does not. For the small-scale qft_6 kernel on the 2D square, Canopus also achieves the optimal routing, superior to TOQM in both $\mathrm{Can}$ and 2Q depth. For the large-scale qft_12 kernel, Canopus consistently outperforms TOQM in both metrics.
@@
Figure 9. QFT kernel fidelity comparison benchmarked on IBM® Quantum Platform (ibm_marrakesh). ibm_marrakesh is the Heron-series QPU with native gate set $\bigl\{\mathrm{CZ},\,\sqrt{\mathrm{X}},\,\mathrm{Z}(\theta),\mathrm{ZZ}(\theta)\bigr\}$.
@@
To further validate these results, we performed real-machine experiments on IBM’s ibm_marrakesh QPU. We compiled QFT circuits of sizes $n\in\{6,8,10,12\}$ for a 1D chain topology using both Canopus and the default Qiskit compiler. Although ibm_marrakesh has a heavy-hex topology, it contains linear chains of sufficient size for these benchmarks. Fidelity was measured using the Hellinger fidelity between the experimental and ideal output distributions, with the number of shots set to $\textsc{max}\{4096,2^{n}\times 10\}$. A layer of Hadamard gates is appended to each circuit execution so that the ideal final state will be $\lvert 0\rangle^{\otimes n}$. In Figure 9, circuits compiled with Canopus achieve, on average, a 52.9% reduction in $\mathrm{CZ}$ gate count, a 66.4% reduction in 2Q-gate depth, and a 26.89% error reduction for the $\mathrm{CZ}/\mathrm{CX}$ and 34.98% for the $\mathrm{ZZ}(\theta)$ gate set, respectively, compared to Qiskit with default settings. These results unequivocally demonstrate the practical advantages of Canopus for QFT kernel compilation.
5.2. qLDPC stabilizer measurement
Figure 10. Logical error rates with error correction via QLDPC stabilizer circuits compiled for 2D heavy-hex (left) and square (right) topologies.
@@
For our another case study, we shift to the fault-tolerant quantum computing (FTQC) context by looking at a kind of important quantum error correction circuit—the stabilizer measurement circuit for qLDPC codes. qLDPC codes are rapidly moving from a topic of theoretical interest to a cornerstone of experimental FTQC research, mainly because of their superior encoding efficiency (bravyi2024high; breuckmann2021quantum). However, due to their frequent long-range interactions for stabilizer measurement (breuckmann2021quantum; panteleev2021degenerate), realizing qLDPC codes on superconducting processors with fixed, local connectivity is still hampered by significant routing overheads (wang2025demonstration).
We demonstrate that the ISA-aware optimization mechanism of Canopus is crucial to mitigating the routing overhead across a diverse set of qLDPC codes. Here we attempt to compile the stabilizer measurement circuits with two ISAs: (1) CX ISA with $\mathrm{CX}$ as the 2Q basis gate; (2) Stab ISA with both $\mathrm{CX}$ and $\mathrm{iSWAP}$ as basis gates, assumed to have an identical cost. Particularly, the Stab ISA aligns with practical hardware realities, e.g., both $\mathrm{CZ}$ and $\mathrm{iSWAP}$ can be natively supported by mainstream superconducting platforms (krantz2019quantum; wei2024native; arute2019quantum). In addition, an ISA incorporating both $\mathrm{iSWAP}$ and $\mathrm{CX}$ leads to significant opportunities to “piggyback” a $\mathrm{SWAP}$ insertion on a $\mathrm{CX}$ without incurring extra 2Q gate count, as the composite block is equivalent to an $\mathrm{iSWAP}$, enabling the possibility of optimizing qubit routing overhead during the execution of stabilizer measurements.
We further build an end-to-end evaluation pipeline with qLDPC code examples from (wang2025demonstration; panteleev2021degenerate), including the generalized bicycle (GB) and bivariate bicycle (BB) codes. We simulate the standard memory experiments using stim (Gidney2021StimAF) to evaluate the fault-tolerant performance of our compiled stabilizer measurement circuits, under the same circuit-level noise model as described in (bravyi2024high). Finally, all syndromes are decoded using the BP-OSD decoder (panteleev2021degenerate; hillmann2024localized) to determine the logical qubit error rate.
@@
As shown in Figure 10, Canopus consistently achieves lower logical error rates than Sabre, as the ISA-aware approach of Canopus results in compiled circuits with less $\mathrm{CX}/\mathrm{iSWAP}$ gate count and circuit depth. Under the CX ISA, Canopus yields an average logical error suppression of 49.4% on the square topology and 11.4% on the heavy-hex topology compared to Sabre. The advantage becomes even more pronounced with the Stab combinatorial ISA, where Canopus achieves a 52.6% (square) and 29.3% (heavy-hex) error suppression, resulting from that there are many opportunities for $\mathrm{SWAP}$ insertions piggybacked on $\mathrm{CX}$ gates without incurring extra 2Q gate count. These results highlight two key findings: first, the ISA-aware mechanism in Canopus is highly effective for compiling QEC circuits, and second, the dedicated use of a hybrid $\mathrm{CX}$-$\mathrm{iSWAP}$ gate set offers a significant practical advantage for qLDPC code demonstrations on superconducting hardware.
@@ 6.1.1. ISAs and basis gate costs
We consider six different ISAs (including the conventional CX ISA) listed in Table 2. These cover a wide range of basis gates from individual $\mathrm{CX}$-family or $\mathrm{iSWAP}$-family gates to combinatorial ones. Particularly, SQiSW (huang2023quantum) proves to a powerful ISA option and has been adopted by recent software projects (mckinney2024mirage; cirqSQiSWDecomposer). ZZPhase ISA containing three fractional $\mathrm{ZZ}(\theta)$ rotation gates (equivalently, $\bigl\{\sqrt[3]{\mathrm{CX}},\sqrt{\mathrm{CX}},\mathrm{CX}\bigr\}$) is adopted by Qiskit’s latest synthesis functionalities (peterson2022optimal; qiskitXXDecomposer). For ZZPhase and SQiSW, we also consider the mirror-enhanced version by incorporating the mirrored basis gates (cross2019validating; mckinney2024mirage) into the ISAs. We also include the Het ISA that is the composition of ZZPhase and SQiSW.
@@
(6) $\left\{\begin{array}[]{c}\mathrm{CX}:1,\,\mathrm{ZZ}(\frac{\pi}{t}):\frac{2}{t},\,\mathrm{\sqrt{iSWAP}}:0.75,\\ \mathrm{iSWAP}:1.5,\,\mathrm{ECP}:1.25,\,\mathrm{pSWAP}(\frac{\pi}{t}):2-\frac{1}{t}\end{array}\right\},$
where $\mathrm{CX}$ gate is the unit cost. Such a setting ensures the continuity of gate costs along the critical edges in the Weyl chamber. For example, $\mathrm{pSWAP}(\pi/2)$ is equivalent to $\mathrm{iSWAP}$ and they have the same cost of $1.5$. With a specific gate family, basis gates with larger canonical coefficients usually requires proportionally longer interaction time on physical devices, which was reflected in the cost setting. Note that this setting is a comprehensive consideration for current gate schemes and hardware-implemented gate fidelities in superconducting (chen2025efficient; arute2019quantum; wei2024native; nguyen2024programmable; acharya2024quantum). It is neither limited to a specific gate scheme nor a specific hardware platform.
@@ 6.1.4. Baselines
Hyperparameters for Sabre and Canopus are of the same settings. Each performs 10 times layout procedure, within which 5-round bidirectional passes are proceeded and each pass performs 10 trials. The best result across all attempts is selected. TOQM can obtain the deterministic routing result in one go. Compiled circuits by BQSKit, although in terms of only the 2Q gate arrangement, is also random. Thus we perform 3 trials for each input case and report the best result.
@@ 6.2. Suppression of routing overhead
Table 4. Average (geometric-mean) routing overhead.
@@
The comprehensive evaluation (on geometric-mean average) results are illustrated in Figure 11 with the average effects summarized in Table 4, across the six selected ISAs and three hardware topologies (1D chain, 2D heavy-hex, 2D square), covering a total of 216 cases ($3\times 6\times 12$) for each compiler. Canopus results in the lowest average (geometric-mean) routing overhead for each ISA-topology combination benchmarking, significantly surpassing other methods. Specifically, the routing overhead reduction achieved by Canopus is on average 16.06% (26.44%), 34.70% (21.25%), and 19.89% (20.72%) in terms of $C_{\mathrm{count}}$ ($C_{\mathrm{depth}}$), compared to Sabre, TOQM, and BQSKit, respectively. With Canopus, those more powerful ISAs, such as ZZPhase and SQiSW, actually exhibit significantly lower routing overhead than CX. While the baseline methods do not apparently indicate that, for which the circuit cost reduction stems from ISA rebase itself.
Notably, Sabre and BQSKit focus on gate count driven optimization, TOQM specializes in lowering circuit depth, while our Canopus involves both count and depth related optimization. Among them, TOQM leads to the worst count-related routing overhead. Even for the routing overhead in terms of depth, TOQM consistently underperforms Canopus, and it only outperforms BQSKit for 2D square topology. In addition, Canopus maintains routing overhead—in terms of both count and depth costs—at a consistently low level; while every baseline fails to optimize some specific circuits. For example, TOQM/BQSKit cannot manage the routing overhead for qec9 circuit; BQSKit cannot for bv circuit, even though with more expressive ISAs.